There are three consecutive intergers listed in natural order. The sum of the first and half the third is 13 less than the second. (What is natural order?)
Thanks

Answer by ReadingBoosters(3246)

x x+1 x+2
x + 1/2(x+2) = x+1 - 13
Combine like terms
x + 1/2x + 1/2(2) = x - 12
1 1/2x + 1 = x - 12
1 1/2x - x = -12 - 1
1/2x = -13
x = -26
Integers are -26, -27 and -28
Proof
(-26)+(1/2)(-28) = -26 + -14 = -40
-27 - 13 = -40